OpenStudy (anonymous):

one solution of the equation (2x-9) (5x+2)=0 this problem is driving me crazy!!!

OpenStudy (anonymous):

x = 9/2 or x = -2/5

razor99 (razor99):

thats the answer

OpenStudy (anonymous):

can you work the problem for me

OpenStudy (anonymous):

see it is product of two factors and their product = 0 this means that one of them has to be = 0 if 2x - 9 = 0 => 2x = 9 => x = 9/2 if 5x + 2 = 0 => 5x = -2 => x = -2/5

razor99 (razor99):

good work @Harkirat

OpenStudy (anonymous):

ohh the question is multiple choice but i was trying to plug the numbers in and cudnt get the answer like that

razor99 (razor99):

lol

OpenStudy (anonymous):

can u get the answer by pluggin the numbers in?

razor99 (razor99):

what do u mean plugging

OpenStudy (anonymous):

Yes.....u can substitute for x ...but one value at a time....